Fermi National Laboratory


Accelerator Update

Friday April 9
The day shift began with Operations monitoring store 3363, with the Antiproton Source (Pbar) not stacking due to a Debuncher quadrupole (D:QS926) problem, with the Recycler maintaining its stash, and with Switchyard (SY) and MiniBooNE taking beam.

Technicians replaced the D:QS926 shunt electronics at 8:13 AM.

At 9:11 AM, the TeV suffered a sector B0 quench.

Pbar resumed stacking at 9:16 AM.

At 10:27 AM, a RF tech finished replacing the bias supply for a Main Injector RF station (MIRF7).

Operations turned off beam to the Meson MCenter experiment at 11:24 AM for an access.

A TeV expert began a twenty-minute study at 12:18 PM.

Operations put the TeV into shot setup around 12:40 PM.

Despite having collimator problems, Operations established store 3366 at 3:12 PM with an initial luminosity of 44.72E30. Pbar resumed stacking.

At 4:46 PM, Operations turned off MIRF to clean the pond water strainer at MI-60. FESS personnel completed this work by 5:21 PM.

Operations set up the Recycler to stash antiprotons from Pbar at 5:41 PM. Pbar resumes stacking at 7 PM.

Operations began sending beam to the Meson MTest and MCenter fixed target experiments at 7:11 PM. MTest turned off for the night at 11 PM.

Saturday April 10
The midnight shift began with Operations monitoring store 3366 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

A MI expert began a coalescing study that began at 9:28 AM and ended at 10:46 AM.

A NuMI expert began a study period that began at noon and finished at 2 PM.

At 2:14 PM, Operations terminated store 3366.

Booster suffered a vacuum burst, tripping off two RF stations (BRF7 and 8). The stations reset okay.

Operations put the TeV through a dry squeeze at 2:28 PM and then into shot setup at 2:50 PM.

BRF13 and 14 tripped off due to a vacuum burst 3:16 PM. They both reset okay.

At 4:52 PM, Operations established store 3368 with an initial luminosity of 48.27E30.

Recycler began stashing antiprotons at 7:57 PM.

Pbar resumed stacking at 8:41 PM.

At 11:32 PM, Linac tripped off due to a vacuum problem with the Linac Buncher. Operators tried to reset the systems, but it didn't take. It seemed that the vacuum went bad, but the pumps remained on. Operators watched as the vacuum came back into its normal range and reset Linac at 11:51 PM.

Sunday April 11
The midnight shift began with Operations monitoring store 3368 and the Recycler stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

MCenter went into a three-hour access around 4 PM.

At 8:56 PM, Operations terminated store 3368. Operators made the Collider Detector at Fermilab (CDF) and D-zero (D0) collision halls ready for access. The accesses were finished by 10 PM.

Monday April 12
The midnight shift began Operations monitoring the Recycler stash, with MiniBooNE taking beam, but with MCenter beam on hold until the new store gets established.

Operations put the TeV into shot setup at 1:24 AM and at 3:11 AM established store 3371 with an initial luminosity of 44.92E30. Pbar resumed stacking and MCenter resumed taking beam.

CDF reported at 3:19 AM a large Pbar halo. Operators investigated and found that some of the collimators didn't move to their correct positions. Neither experiment could use the beam for data taking due to this problem. Operators contacted experts for help.

At 4:11 AM, the TeV suffered a sector F4 quench due to a collimator moving to a bad position.

The TeV turn-on began around 6:30 AM.

Plans for Today
The plan for today is to allow TeV experts to conduct studies. Later on today the Recycler will stash antiprotons

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/12/2004   email Fermilab